Shift summary. Eve 08/01/99

Patrick Toale (toale@cuhep.Colorado.EDU)
Sun, 1 Aug 1999 23:03:14 -0600 (MDT)

Shift Summary for Eve shift, 08/01/99

On Shift: Pat T., Sydney T.

Hours of beam delivered at all intensities: 8.0/8.5
Hours of beam delivered at nominal intensity: 8.0/8.5
Hours of E832 data (special runs not included): 5.0/8.5
Hours of special runs (laser scans, muon runs...): 3.5/8.5

Accounting of lost time: new tapes, naughty chipmunk in booster,
swapping bad dc-or power supply

========================================================================
Runs ended this shift: 14078, 14080, 14081, 14083
Run 14078 - 24.4M events - BEAM_832
Run 14080 - 2.3M events - BEAM_832
Run 14081 - COSMICS
RUN 14083 - PEDS

Tapes brought to FCC this shift: QK*204-206, QKB070, QKA005, QKCC08,
QKLA10, QKN016-017, QKNL15-16, QKSM08
(including tapes for online split)

========================================================================
Accesses during shift: Hogan N. and Sydney T. make access to replace
bad dc-or power supply on DC2Y.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14078 15:30 Inherit smooth run

16:50 DN puts 2 new argon-ethane bottles on line on
east side

19:05 Tapes full. Stop run 14078

19:20 Run 14079 fails to start: "DA_START FAILED" pink
window appears => drt_fresh_start

14080 19:27 Start BEAM_832 run 14080

19:30 Notice BA_SUM pedestals off by ~20 counts,
Elizabeth paged

19:45 Elizabeth arrives, BA_SUM pedestals are only
off for pre-bucket sums which are not really used so
continued running with current peds

31 20:11 Beam goes away, bad chipmunk in booster, est 15 min.

46 20:27 Beam back

57 20:45 DC2Y starts to oscillate, sdaq reports DC2STORY_2
is putting out -0.7V instead of nominal -5.0V,
Hogan paged

21:05 Stop run 14080

14081 21:12 Start COSMICS run 14081

21:20 HN and ST go on access to measure output of DC2Y
stretcher PS, find it to be -1.8, page Ashkan, call
Theo

23:30 HN and ST out of access after replacing DC2Y
stretcher-OR power supply

23:35 Stop run 14081

14083 23:40 Start PED run 14083

23:48 Run 14083 stopped

14084 23:52 Start BEAM_832 run 14084
ET Top-West quadrant drops out for first
3 spills. Reload ET.

24:00 Hand-off to owl shift.

========================================================================
Follow up for next shift:

Thanks,
Pat Toale
--------------------------------------------------------------
Graduate Research Assistant
Department of Physics University of Colorado at Boulder
Office: F-231 Phone: (303) 492-4363
E-Mail: toale@cuhep.colorado.edu